Two cards are dealt in succession and without replacement from a standard deck of 52 cards.
(a) In how many different orders can the cards be dealt?
(b) How many different 2-card hands are possible?

Answer by edjones(8007)

(a) In how many different orders can the cards be dealt?
Permutation: nPr=52P2=52!/(52-2)!=52*51=2652 order is important.
Combination: nCr=52C2=52!/((52-2)!2!)=1326 order is not important.
.
Ed